# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.4 Motivation
Finally, note that if we had chosen to take time steps of length $\operatorname{d}t$ instead of $\sqrt{\operatorname{d}t}$ under pushforwards of $\rho^{X-\overline{X}}$, the first term of the RHS in Definition 3.9 would have become $\int_{M}\int_{\mathcal{X}(M)}\mathcal{L}_{X-\overline{X}_{V(\mu(t))}}(\varphi)\operatorname{d}V(\mu(t))\operatorname{d}\mu(t)$. Hence, Definition 3.9 would become
$\frac{\operatorname{d}}{\operatorname{d}t}\int_{M}\varphi\operatorname{d}\mu(t)=\int_{M}\int_{\mathcal{X}(M)}\mathcal{L}_{X}\varphi\operatorname{d}V(\mu(t))\operatorname{d}\mu(t),$ (3.39)
@@ 3.4.1 Examples
If we take $M=\mathbb{R}$ and $V(\mu)=\eta$ for all $\mu\in\mathcal{P}(M)$, where $\eta$ gives weight $\frac{1}{2}$ to the vector fields $X_{1}$ and $X_{-1}$ that are constant $1$ and 1-1, respectively, then the Wiener process, i.e., $\mu(t)=\frac{1}{\sqrt{2\pi t}}e^{-\frac{x^{2}}{2t}}\operatorname{d}x$ is a solution of (MDE) with $\mu_{0}=\delta_{0}$. This is a good sign, as the Wiener process is obtained as a limit of random walks with step-size going to 0.
Analogously if we set $\eta$ to be the measure that gives weight $\frac{1}{2}$ to the vector fields $X_{1}$ and $X_{-1}$, then for any $T$, the backward Wiener process with $\mu(t)=\delta_{0}$ for $t>T$ is a solution of (MDE) with $\mu_{0}=f_{T}(x)dx$ with $f_{t}:x\mapsto\frac{1}{\sqrt{2\pi t}}\exp{-\frac{x^{2}}{2t}}$.
@@
Let us now give a more sophisticated example by choosing $M=\mathbb{R}^{2}$. Otherwise we stick to the notations of the previous examples, this time giving equal weights of $\frac{1}{3}$ to the roots of $X^{3}-1$, which we denote by $j_{1}$, $j_{2}$ and $j_{3}$. For $\mu_{0}=\delta_{0}$ we expect to get some Gaussian diffusion with a rotational symmetry, i.e., whose covariance-matrix is a multiple of the identity matrix. Indeed, in this case one readily checks that the operator $\frac{1}{3}(\frac{\partial^{2}}{\partial j_{1}^{2}}+\frac{\partial^{2}}{\partial j_{2}^{2}}+\frac{\partial^{2}}{\partial j_{3}^{2}})=\frac{1}{2}\Delta$, where $\Delta$ denotes the Laplace operator—which gives a solution of the desired form. The detailed computation goes
@@ 3.5 The differential operator $\square_{\mu}^{V}$
Let us now study some basic properties of the differential operator $\square_{\mu}^{V}$. First, we now make a stronger assumption on $V$, namely $(\pi_{p})_{*}V[\mu]\in\mathcal{W}_{2}(M)$ for all $p\in M$ and $\mu\in\mathcal{P}(M)$ — or directly $V[\mu]\in\mathcal{W}_{2}(\mathcal{X}^{2,\infty})$. Then $(\pi_{p})_{*}V[\mu]$ is in particular an integrable signed Borel–measure. Let $\nu_{+}$ and $\nu_{-}$ be the unique pair of positive 22–integrable Borel–measures such that $(\pi_{p})_{*}V[\mu]=\nu_{+}-\nu_{-}$. Furthermore, choose a smooth frame $(X_{1},\dots,X_{n})$ on $M$ and take $X\in\mathcal{X}(M)$. Then for $X=X_{i}\frac{\partial}{\partial x_{i}}\in\mathcal{X}(M)$, in Einstein–convention,
@@
We are particularly interested in the term of second order of $\square_{\mu(t)}^{V}$. Note that the second order term $X_{i}X_{j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}$ is tensorial in $X$, so it suffices to study it pointwise on tangent spaces. Since $(\pi_{p})_{*}V[\mu]\in\mathcal{W}_{2}(M)$, for all $\varphi\in\mathcal{C}^{\infty}_{c}(M)$, $\int_{T_{p}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(p)\operatorname{d}(\pi_{p})_{*}V[\mu](X)$ converges, and
$\int_{T_{p}M}$ $X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(p)\operatorname{d}(\pi_{p})_{*}V[\mu](X)$ (3.43)
$=\int_{T_{p}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(p)\operatorname{d}\nu_{+}(X)$
$\qquad-\int_{T_{p}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(p)\operatorname{d}\nu_{-}(X)$
$=\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(p)\biggl{(}\int_{T_{p}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)-\int_{T_{p}M}X_{i}X_{j}\operatorname{d}\nu_{-}(X)\biggr{)},$
where $\pi_{i}$ denotes the projection on $x_{i}$. Indeed, for all $i,j$ and all $X\in T_{p}M$, $X_{i}X_{j}\leq\|X\|_{2}^{2}$ and by assumption $X\mapsto\|X\|_{2}^{2}$ is $\nu_{+}$– and $\nu_{-}$–integrable. Indeed, $\int_{T_{p}M}\|X\|_{2}^{2}\operatorname{d}\nu_{+}(X)$ is just the $2$-moment of $\nu_{+}(X)$. Now, for all $i,j$, $\int_{T_{p}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)=\langle\pi_{i},\pi_{j}\rangle_{L^{2}_{\nu_{+}}}$. Hence, $\bigl{(}\int_{T_{p}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)\bigr{)}_{i,j}$ is a Gram–matrix and, thus, symmetric and semi-elliptic. With the same reasoning, we see that $\bigl{(}\int_{T_{p}M}X_{i}X_{j}\operatorname{d}\nu_{-}(X)\bigr{)}_{i,j}$ is symmetric and semi-elliptic. This yields a decomposition of $\square_{\mu}^{V}=\Delta_{\mu(t)}^{+,V}-\Delta_{\mu(t)}^{-,V}$ into two point–wise semi-elliptic operators, i.e., equation (MDE) is of the form
@@
If $V$ takes values in the positive Borel–measures or, in particular, some $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$, $\Delta_{\mu}^{-,V}$ vanishes for all $\mu$ and equation (3.14) resembles a semi-parabolic equation. We shall see that, in that case, the assumptions on $\square_{\mu}^{V}$ that are necessary to guarantee the existence of a solution of (3.14) are very similar to those we know from the classical solution theory of semi-parabolic PDEs.
@@
$a_{i,j}^{v}(x)$ $\coloneqq\int_{T_{p}M}X_{i}(x)X_{j}(x)\operatorname{d}v(X)$ (3.45)
$b_{j}^{v}(x)$ $\coloneqq\int_{T_{p}M}X_{i}(x)\frac{\partial X_{j}}{\partial x_{i}}(x)\operatorname{d}v(X).$
@@
Let $p\geq 2$, $V$ be a vector–field probability, such that there exists $B>0$, such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])\leq B$. Let $\varepsilon>0$. Then, there exists a vector–field probability $W$ such that $\mathcal{W}_{p}^{\infty}(V,W)\leq\varepsilon$ and $\square^{W}_{\mu}$ is $\mathcal{K}\varepsilon$–elliptic for all $\mu\in\mathcal{W}_{p}(M)$, and $\mathcal{K}>0$ depends only on $B$ and $M$.
@@
where $\pi_{i(T_{x}M)}$ denotes the orthogonal projection on the space $i(T_{x}M)$. Then, $\bigl{(}X_{j}\bigr{)}_{j\in\llbracket 0,N-1\rrbracket}$ is a family of smooth vector–fields and $\bigl{(}X_{j}(x)\bigr{)}_{j\in\llbracket 0,N-1\rrbracket}$ is a generating system of $T_{x}M$ for all $x\in M$. By construction, for all $j\in\llbracket 0,N-1\rrbracket$ and all $x\in M$, $\|X_{j}(x)\|\leq 1$.
@@
$\qquad\qquad\qquad\qquad+\mathcal{W}_{p}((1-\varepsilon)V[\mu]+\varepsilon\delta_{0},W[\mu])$
$\leq(B+1)\varepsilon.$
Finally, it follows immediately from (3.43) that $\square^{W}_{\mu}$ is $\frac{\varepsilon}{N}$–elliptic for all $\mu\in\mathcal{W}_{p}(M)$. $\square$
3.6 Existence and construction of solutions
We shall now prove the existence of solutions of (3.14) by a discrete–to–continuous argument, as sketched in Section 3.4. This proof requires that $V[\mu]$ has finite $p$-th moment for some fixed $p>2$ and for all $\mu\in\mathcal{W}_{p}(M)$ with respect to the $\|\leavevmode\nobreak\ \|_{W^{2,\infty}}$–norm. Therefore, we have to make a stronger assumption than before, namely $p>2$.
@@ 3.6 Existence and construction of solutions
Let $T>0$. Let $V$ be a uniformly continuous probability vector field on $\mathcal{W}_{p}(M)$ for $p>2$. Assume that $V$ has uniformly bounded pp–th moments, e.g, there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume furthermore that $(M,g)$ is of bounded geometry and that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Then, (3.14) has a $\frac{1}{p}$–Hölder–continuous solution in $\mathcal{W}_{p}(M)$ on $[0,T]$.
@@
    Assume that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, (3.14) has a continuous solution in $\mathcal{W}_{p}(M)$ on $[0,T]$.
We will prove Theorem 3.26 as a Corollary of the stronger Theorem 3.35, which is in some sense a generalisation of the well-known Central Limit Theorem. It affirms the existence of solutions to (3.14) in a constructive way.
@@ 3.6.1 Average Flow Approximation Series (AFAS)
An explicit solution of (3.14) is obtained as a limit of an explicit Euler–scheme—which we will call Average Flow Approximation Series (AFAS).
@@
for all $t\in[\frac{x_{l+1}+x_{l}}{2},x_{l}]$ and
$\mu^{\mathcal{}\mathfrak{P}(\delta)}(t)=\bigl{(}e^{V}_{2(t-\frac{x_{l+1}+x_{l}}{2})}\circ f^{V}_{x_{l+1}-x_{l}}\bigr{)}\bigl{(}\mu^{\mathcal{}\mathfrak{P}(\delta)}\bigl{(}\frac{x_{l+1}+x_{l}}{2}\bigr{)}\bigr{)},$
for all $t\in[x_{l+1},\frac{x_{l+1}+x_{l}}{2}]$.
As proposed in Section 3.4, we call a solution of (3.14) smooth if it is a limit of AFAS.
@@
Let $p\geq 2$ and $T>0$. A narrowly continuous curve $\nu:[0,T]\to\mathcal{W}_{p}(M)$ is called a smooth solution of (3.14) if it solves (3.14) and there exists a sequence $(\mathcal{}\mathfrak{P}_{n})_{n\in\mathbb{N}}$ of partitions of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}_{n})\to 0$ such that for all $n\in\mathbb{N}$, $\mu^{\mathcal{}\mathfrak{P}_{n}}$ converges pointwise narrowly to $\nu$.

### v2

@@ 3.4 Motivation
Finally, note that if we had chosen to take time steps of length $\operatorname{d}t$ instead of $\sqrt{\operatorname{d}t}$ under pushforwards of $\rho^{\overline{X}[\mu(t)]}$, the first term of the RHS in Definition 3.11 would have become $\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\mathcal{L}_{\overline{X}[\mu(t)]}(\varphi)\operatorname{d}V[\mu(t)]\operatorname{d}\mu(t)$. Hence, Definition 3.11 would become
$\frac{\operatorname{d}}{\operatorname{d}t}\int_{M}\varphi\operatorname{d}\mu(t)=\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\mathcal{L}_{X}\varphi\operatorname{d}V[\mu(t)](X)\operatorname{d}\mu(t),$ (3.38)
@@ 3.4.1 Examples
If we take $M=\mathbb{R}$ and $V(\mu)=\eta$ for all $\mu\in\mathcal{P}(M)$, where $\eta$ gives weight $\frac{1}{2}$ to the vector–fields $X_{1}$ and $X_{-1}$ that are constant $1$ and 1-1, respectively, then the Wiener process, i.e., $\mu(t)=\frac{1}{\sqrt{2\pi t}}e^{-\frac{x^{2}}{2t}}\operatorname{d}x$ is a solution of (MDE) with $\mu_{0}=\delta_{0}$. This is a good sign, as the Wiener process is obtained as a limit of random walks with step-size going to 0.
Analogously if we set $\eta$ to be the measure that gives weight $\frac{1}{2}$ to the vector–fields $X_{1}$ and $X_{-1}$, then for any $T$, the backward Wiener process with $\mu(t)=\delta_{0}$ for $t>T$ is a solution of (MDE) with $\mu_{0}=f_{T}(x)dx$ with $f_{t}:x\mapsto\frac{1}{\sqrt{2\pi t}}\exp{-\frac{x^{2}}{2t}}$.
@@
Let us now give a more sophisticated example by choosing $M=\mathbb{R}^{2}$. Otherwise we stick to the notation of the previous examples, this time giving equal weights of $\frac{1}{3}$ to the roots of $X^{3}-1$, which we denote by $j_{1}$, $j_{2}$ and $j_{3}$. For $\mu_{0}=\delta_{0}$ we expect to get some Gaussian diffusion with a rotational symmetry, i.e., whose covariance-matrix is a multiple of the identity matrix. Indeed, in this case one readily checks that the operator $\frac{1}{3}(\frac{\partial^{2}}{\partial j_{1}^{2}}+\frac{\partial^{2}}{\partial j_{2}^{2}}+\frac{\partial^{2}}{\partial j_{3}^{2}})=\frac{1}{2}\Delta$, where $\Delta$ denotes the Laplace operator—which gives a solution of the desired form. The detailed computation goes
@@ 3.5 The differential operator $\square_{\mu}^{V}$
Let us now study some basic properties of the differential operator $\square_{\mu}^{V}$. First, we now make a stronger assumption on $V$, namely $(\pi_{x})_{*}V[\mu]\in\mathcal{W}_{2}(M)$ for all $x\in M$ and $\mu\in\mathcal{P}(M)$ — or directly $V[\mu]\in\mathcal{W}_{2}(\mathcal{X}^{2,\infty}(M))$. Then $(\pi_{x})_{*}V[\mu]$ is in particular an integrable signed Borel–measure. Let $\nu_{+}$ and $\nu_{-}$ be the unique pair of positive 22–integrable Borel–measures such that $(\pi_{x})_{*}V[\mu]=\nu_{+}-\nu_{-}$. Furthermore, choose a smooth frame $(X_{1},\dots,X_{n})$ on $M$ and take $X\in\mathcal{X}^{2,\infty}(M)$. Then for $X=X_{i}\frac{\partial}{\partial x_{i}}\in\mathcal{X}^{2,\infty}(M)$, in Einstein–convention,
@@
We are particularly interested in the term of second order of $\square_{\mu(t)}^{V}$. Note that the second order term $X_{i}X_{j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}$ is tensorial in $X$, so it suffices to study it pointwise on tangent spaces. Since $(\pi_{x})_{*}V[\mu]\in\mathcal{W}_{2}(M)$, for all $\varphi\in\mathcal{C}^{\infty}_{c}(M)$, $\int_{T_{x}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(x)\operatorname{d}(\pi_{x})_{*}V[\mu](X)$ converges, and
$\int_{T_{x}M}$ $X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(x)\operatorname{d}(\pi_{x})_{*}V[\mu](X)$ (3.42)
$=\int_{T_{x}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(x)\operatorname{d}\nu_{+}(X)$
$\hskip 18.49988pt-\int_{T_{x}M}X_{i}X_{j}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(x)\operatorname{d}\nu_{-}(X)$
$=\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}(x)\biggl(\int_{T_{x}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)-\int_{T_{x}M}X_{i}X_{j}\operatorname{d}\nu_{-}(X)\biggr),$
where $\pi_{i}$ denotes the projection on $x_{i}$. Indeed, for all $i,j$ and all $X\in T_{x}M$, $X_{i}X_{j}\leq\|X\|_{2}^{2}$ and by assumption $X\mapsto\|X\|_{2}^{2}$ is $\nu_{+}$– and $\nu_{-}$–integrable. Indeed, $\int_{T_{x}M}\|X\|_{2}^{2}\operatorname{d}\nu_{+}(X)$ is just the $2$-moment of $\nu_{+}(X)$. Now, for all $i,j$, $\int_{T_{x}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)=\langle\pi_{i},\pi_{j}\rangle_{L^{2}_{\nu_{+}}}$. Hence, $\bigl(\int_{T_{x}M}X_{i}X_{j}\operatorname{d}\nu_{+}(X)\bigr)_{i,j}$ is a Gram–matrix and, thus, symmetric and semi-elliptic. With the same reasoning, we see that $\bigl(\int_{T_{x}M}X_{i}X_{j}\operatorname{d}\nu_{-}(X)\bigr)_{i,j}$ is symmetric and semi-elliptic. This yields a decomposition of $\square_{\mu}^{V}=\Delta_{\mu(t)}^{+,V}-\Delta_{\mu(t)}^{-,V}$ into two point–wise semi-elliptic operators, i.e., equation (MDE) is of the form
@@
If $V$ takes values in the positive Borel–measures or, in particular, some $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$, $\Delta_{\mu}^{-,V}$ vanishes for all $\mu$ and equation (MDEd) resembles a semi-parabolic equation. We shall see that, in that case, the assumptions on $\square_{\mu}^{V}$ that are necessary to guarantee the existence of a solution of (MDEd) are very similar to those we know from the classical solution theory of semi-parabolic PDEs.
@@
$a_{i,j}^{v}(x)$ $\coloneqq\int_{T_{x}M}X_{i}(x)X_{j}(x)\operatorname{d}v(X)$ (3.44)
$b_{j}^{v}(x)$ $\coloneqq\int_{T_{x}M}X_{i}(x)\frac{\partial X_{j}}{\partial x_{i}}(x)\operatorname{d}v(X).$
@@
Let $p\geq 2$, $(M,g)$ be a Riemannian manifold of bounded geometry and $V$ be a probability vector–field, such that there exists $B>0$, such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])\leq B$. Let $\varepsilon>0$. Then, there exists a probability vector–field $W$ such that
  • $\mathcal{W}_{p}^{\infty}(V,W)\leq\varepsilon^{\frac{1}{p}}$.
  • $\square^{W}_{\mu}$ is $\mathcal{K}\varepsilon$–elliptic for all $\mu\in\mathcal{W}_{p}(M)$, for a constant $\mathcal{K}>0$ depending only on $B$ and $M$.
  • for all $\mu,\nu\in\mathcal{W}_{p}(M)$, $\mathcal{W}_{p}\bigl(W[\mu],W[\nu]\bigr)\leq\mathcal{W}_{p}\bigl(V[\mu],V[\nu]\bigr)$.
  • $W$ has uniformly bounded pp–th moments.
@@
where $\pi_{i(T_{x}M)}$ denotes the orthogonal projection on the space $i(T_{x}M)$. Then, $\bigl(X_{j}\bigr)_{j\in\llbracket 0,N-1\rrbracket}$ is a family of smooth vector–fields and $\bigl(X_{j}(x)\bigr)_{j\in\llbracket 0,N-1\rrbracket}$ is a generating system of $T_{x}M$ for all $x\in M$. By construction, for all $j\in\llbracket 0,N-1\rrbracket$ and all $x\in M$, $\|X_{j}(x)\|\leq 1$. Furthermore, since $(M,g)$ is of bounded geoemtry, there exists $C(M)>0$, such that for all $j\in\llbracket 0,N-1\rrbracket$, $\|X_{j}\|_{W^{2,\infty}}\leq C(M)$.
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt+\mathcal{W}_{p}((1-\varepsilon)V[\mu]+\varepsilon\delta_{0},W[\mu])$
$\leq(B^{\frac{1}{p}}+C(M))\varepsilon^{\frac{1}{p}}.$
Finally, it follows immediately from (3.42) that $\square^{W}_{\mu}$ is $\frac{\varepsilon}{N}$–elliptic for all $\mu\in\mathcal{W}_{p}(M)$.
The last two points are immediate. $\square$
3.6 Existence and construction of solutions of (MDE)
We shall now prove the existence of solutions of (MDEd) by a discrete–to–continuous argument, as sketched in Section 3.4. This proof requires that $V[\mu]$ has finite $p$-th moment for some fixed $p>2$ and for all $\mu\in\mathcal{W}_{p}(M)$ with respect to the $\|\penalty 10000\ \|_{W^{2,\infty}}$–norm. Therefore, we have to make a stronger assumption than before, namely $p>2$.
@@ 3.6 Existence and construction of solutions of (MDE)
Let $T>0$. Let $V$ be a uniformly continuous probability vector–field on $\mathcal{W}_{p}(M)$ for $p>2$. Assume that $V$ has uniformly bounded pp–th moments, e.g., there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume furthermore that $(M,g)$ is of bounded geometry and that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(M)$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Then, (MDEd) has a $\frac{1}{p}$–Hölder–continuous solution in $\mathcal{W}_{p}(M)$ on $[0,T]$.
@@
    Assume only that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, (MDEd) has a continuous solution in $\mathcal{W}_{p}(M)$ on $[0,T]$.
We will prove Theorem 3.28 as a corollary of the stronger Theorem 3.39, which is in some sense a generalisation of the well-known Central Limit Theorem. It affirms the existence of solutions to (MDEd) in a constructive way.
@@ 3.6.1 Average Flow Approximation Series (AFAS)
An explicit solution of (MDEd) is obtained as a limit of an explicit Euler–scheme—which we will call Average Flow Approximation Series (AFAS).
@@
for all $t\in\bigl[x_{l},\frac{x_{l+1}+x_{l}}{2}\bigr]$ and
$\mu^{\mathcal{}\mathfrak{P}(\delta)}(t)=\biggl(e^{V}_{2\bigl(t-\frac{x_{l+1}+x_{l}}{2}\bigr)}\circ f^{V}_{x_{l+1}-x_{l}}\biggr)\bigl(\mu^{\mathcal{}\mathfrak{P}(\delta)}(x_{l})\bigr),$
for all $t\in\bigl[\frac{x_{l+1}+x_{l}}{2},x_{l+1}\bigr]$.
As proposed in Section 3.4, we call a solution of (MDEd) smooth if it is a limit of AFAS.
@@
Let $p\geq 2$ and $T>0$. A narrowly continuous curve $\nu:[0,T]\to\mathcal{W}_{p}(M)$ is called a smooth solution of (MDEd) if it solves (MDEd) and there exists a sequence $(\mathcal{}\mathfrak{P}_{n})_{n\in\mathbb{N}}$ of partitions of $[0,T]$ with maximal step–length $\mathfrak{p}(\mathcal{}\mathfrak{P}_{n})\to 0$ such that for all $n\in\mathbb{N}$, $\mu^{\mathcal{}\mathfrak{P}_{n}}$ converges pointwise narrowly to $\nu$.
3.6.2 Averaged flow operators—Hölder–regularity in time